# Cog-RAG: Cognitive-Inspired Dual-Hypergraph with Theme Alignment Retrieval-Augmented Generation

### v1

@@ Introduction
For a given user query $q$, following (guo2024lightrag; feng2025hyper), we first extract two types of keywords: theme keywords (overarching concepts or themes) and entity keywords (specific entities or details), as follows:
$\mathcal{X_{\text{theme}}},\mathcal{X_{\text{entity}}}=LLM(\mathcal{P_{\text{keyword}}}(q)),$ (9)
where $\mathcal{X_{\ast}}=\{{x_{1},x_{2},...}\}$, $\mathcal{P_{\text{keyword}}}$ is the prompt for extracting keywords from the query, detailed in Appendix.
@@
Subsequently, based on the extracted keywords, we perform structured retrieval over the hypergraph database. It is worth noting that entity keywords primarily describe concrete individual information and are thus matched to vertices, while theme keywords reflect abstract semantic relations among multiple entities and are therefore used to retrieve relevant hyperedges. The natural combination of two types of keywords and hypergraph structure enhances both retrieval specificity and structural compatibility.
In the first stage of retrieval, the extracted theme keywords are used to perform semantic matching within the theme hypergraph, and selecting the top-k relevant theme hyperedges.
@@
where $\mathcal{E}_{\text{rel}}$ represents the relevant hyperedges retrieved from the vector database. Then, we perform a diffusion process over the hypergraph database to retrieve their neighboring vertices, providing additional context awareness of the retrieved theme.
@@
$\mathcal{V}_{\text{rel}}=\{\mathcal{R}(\mathcal{P}_{\text{align}}(x_{i},\mathcal{A}_{\text{theme}}),\mathcal{V}_{\text{entity}})|x_{i}\in\mathcal{X}_{\text{entity}}\},$ (13)
where $\mathcal{P}_{\text{align}}$ is the prompt (detailed in Appendix) used for embedding theme knowledge. $\mathcal{V}_{\text{rel}}$ refers the retrieved relevant entities. Then perform a hypergraph structure diffusion as follows:
@@ F.1 Extracting Themes, Key Entities
Extracting Key Entities Formulation: $\mathcal{P}_{\text{ext\_theme}}$
@@ F.2 Extracting Entities, Relations and Keywords
Extracting Keywords from Query Formulation: $\mathcal{P}_{\text{keywords}}$
Prompt: You are a helpful assistant tasked with identifying both theme-level and entity-level keywords in the user’s query.
@@
Given the query and its related theme, please refer to the theme, list both theme-level and entity-level keywords. theme-level keywords focus on overarching concepts or themes, while entity-level keywords focus on specific entities, details, or concrete terms.
@@
- entities_set: The collection of names for elements in high-order associated entity set. - high_order_relationship_description: Use the relationships among the entities in the set to create a detailed, smooth, and comprehensive description that covers all entities in the set, without leaving out any relevant information.
@@ F.3 Theme Alignment to Entity
Theme Alignment to Entity Formulation: $\mathcal{P}_{\text{algin}}$
Prompt: Through the existing analysis, we can know the potential theme answer and entity-level keywords are $\mathcal{A}_{\text{theme}}$ and $\mathcal{X}_{\text{ext\_entity}}$. Please refer to the theme answer and entity-level keywords, combined with your own analysis, to select useful and relevant information to help you answer accurately.
@@ F.4 Evaluation Metrics
Scoring-based Evaluation Formulation: $\mathcal{P}_{\text{eval\_scoring}}(q,\mathcal{A},\mathcal{T}_{o})$
$q$ denotes user query, $\mathcal{A}$ denotes LLM response, $\mathcal{T}_{o}$ denotes the original text chunk that generated the question.
@@
Here are the relevant document: $\mathcal{T}_{o}$

### v2

@@ Introduction
For a given user query $q$, we first extract theme keywords (overarching concepts or themes), as follows:
$\mathcal{X_{\text{theme}}}=LLM(\mathcal{P_{\text{keyword}}}(q)),$ (9)
where $\mathcal{X_{\ast}}=\{{x_{1},x_{2},...}\}$, $\mathcal{P_{\text{keyword}}}$ is the prompt for extracting theme keywords from the query, detailed in Appendix.
@@
Subsequently, we perform structured retrieval over the hypergraph database. It is worth noting that theme keywords reflect abstract semantic relations among multiple entities and are therefore used to retrieve relevant hyperedges.
Therefore, in the first stage of retrieval, the extracted theme keywords are used to perform semantic matching within the theme hypergraph and select the top-k relevant theme hyperedges.
@@
where $\mathcal{E}_{\text{rel}}$ represents the relevant hyperedges retrieved from the vector database. Then, we perform a diffusion process over the hypergraph database to retrieve their neighboring vertices, providing additional context awareness for the retrieved theme.
@@
Based on the theme response, we further extract the aligned entity keywords (specific entities or details) from query $q$, which is as follows:
$\mathcal{X_{\text{entity}}}=LLM(\mathcal{P_{\text{align}}}(q,\mathcal{A}_{\text{theme}})),$ (13)
where $\mathcal{P}_{\text{align}}$ is the prompt (detailed in Appendix) used for extracting entity keywords aligned with the theme. $\mathcal{X_{\text{entity}}}$ primarily describe concrete individual information and are thus matched to vertices. The natural combination of two types of keywords and hypergraph structure enhances both retrieval specificity and structural compatibility.
@@
$\mathcal{V}_{\text{rel}}=\{\mathcal{R}((x_{i},\mathcal{V}_{\text{entity}})|x_{i}\in\mathcal{X}_{\text{entity}}\},$ (14)
where $\mathcal{V}_{\text{rel}}$ refers the retrieved relevant entities. Then perform a hypergraph structure diffusion as follows:
@@ F.1 Extracting Themes, Key Entities
Extracting Key Entities Formulation: $\mathcal{P}_{\text{ext\_key}}$
@@ F.2 Extracting Entities, Relations and Keywords
Extracting Theme Keywords from Query Formulation: $\mathcal{P}_{\text{keywords}}$
Prompt: You are a helpful assistant tasked with identifying theme-level keywords in the user’s query.
@@
Given the query, list theme-level keywords. theme-level keywords focus on overarching concepts or themes.
@@
- entities_set: The collection of names for elements in high-order associated entity set.
- high_order_relationship_description: Use the relationships among the entities in the set to create a detailed, smooth, and comprehensive description that covers all entities in the set, without leaving out any relevant information.
@@ F.3 Theme Alignment to Entity
Theme Alignment to Entity Formulation: $\mathcal{P}_{\text{algin}}(q,\mathcal{A}_{\text{theme}})$
Prompt: You are a helpful assistant tasked with extracting entity-level keywords from the query $q$ that are theme-aligned.
—Goal—
Given the query and theme, extract specific entities and concrete terms that relate to the identified theme.
- Extract keywords that are concrete and entity-focused.
- Prioritize keywords aligned with the theme.
Here are the question: $q$
Here are the theme: $\mathcal{A}_{\text{theme}}$
@@ F.4 Evaluation Metrics
Scoring-based Evaluation Formulation: $\mathcal{P}_{\text{eval\_scoring}}(q,\mathcal{A},\mathcal{C})$
$q$ denotes user query, $\mathcal{A}$ denotes LLM response, $\mathcal{C}$ denotes the original text chunk that generated the question.
@@
Here are the relevant document: $\mathcal{C}$